OpenStudy (anonymous):

using loarithms solve:the population of a city is growing at a rate proportional to its population. the population 20 years ago was 100 000 and today it is 150 000. find the population 20 years from now

OpenStudy (lilg132):

225,000

OpenStudy (anonymous):

how?

OpenStudy (lilg132):

150,000 / 100,000 = 1.5 150,000 * 1.5 = 225,000

OpenStudy (lilg132):

we divide 150,000 by 100,000 because we want to find out how much the population changed by in that amount of time which is 20 years and we need to find out how much it will change after another 20 years so we use the same percentage

OpenStudy (lilg132):

1.5 = 150% 7.5% every year

OpenStudy (lilg132):

150 / 20 = 7.5 % every year